### v1

@@ 2.2 Evaluation of downscaling and bias correction performance
We investigate the inference performance of our diffusion model on embedded GFDL data $g(GFDL)$. We compare the downscaling and bias correction performance of our diffusion model to a benchmark consisting of first applying bi-linear upsampling followed by QM for bias correction.
Figure 3 allows for a qualitative comparison between the different individual precipitation fields. The upsampled $GFDL$ fields, as well as our benchmark consisting of the upsampled and QM-corrected $GFDL$ fields, are visually too smooth. They therefore appear blurry compared to the ERA5 precipitation fields despite being on the same spatial resolution of 0.25°. Our diffusion model produces detailed, high resolution outputs that are visually indistinguishable from the $ERA5$ reanalysis that we treat as ground truth.
@@
Figure 3: Comparative visualization of individual randomly selected samples: Each row presents four samples of the same dataset. The top row shows GFDL ESM4 data, bi-linearly upsampled to 0.25° to match the other fields. The second row shows QM-corrected and the third row diffusion model-corrected GFDL fields. The bottom row shows samples of the original ERA5 data, which are unpaired to the GFDL fields above. Visual inspection shows that the diffusion model correction greatly improves upon the QM correction in terms of producing realistic spatial patterns, since the QM-corrected fields remain way too blurry compared to the HR ERA5 data. There is no visual difference between the details and sharpness of diffusion model-corrected GFDL fields compared to ERA5.
A comparison between the absolute temporally and absolute spatial-temporally averaged diffusion model-corrected GFDL and ERA5 fields (Fig. 4A) yields a bias of 0.29 mm/d. This is a significant improvement over the original GFDL dataset which yields a bias of 0.69 mm/d (Fig. 4B). Our diffusion model performs on par with the state-of-the-art bias correction performance of our benchmark that is by design optimal for this task, at 0.26 mm/d (Fig. 4D).
@@
Figure 4: Bias of the GFDL ESM4 model and the QM- and diffusion model-corrections, defined as the difference between long-term temporal averages of all validation samples. Specifically, the temporally averaged bias fields with respect to ERA5 are shown for (A) the diffusion model correction, (B) the uncorrected GFDL and (C) the QM correction. Results indicate a substantial improvement of our diffusion model (A) and the benchmark (C) over just upsampling GFDL to 0.25°. The absolute bias on top of each panel is given by the mean absolute value of the differences over the spatial and temporal dimension with respect to ERA5.
@@
Figure 5: Comparison of GFDL (bi-linearly upsampled to 0.25°) (orange) and ERA5 (black) to diffusion model-corrected GFDL (magenta) and QM-corrected GFDL fields (blue) as our benchmark. The Power spectral density (PSD) plot (A) shows that the diffusion model corrects the small-scale spatial details far better then our benchmark. The spectrum aligns very well with the high resolution ERA5 target data. The histograms (B) as well as the latitude (C) and longitude (D) profiles show substantial improvements compared to the uncorrected GFDL data.
@@
There are large differences between GFDL and ERA5 data in small-scale patterns (Fig. 5A). The histogram of precipitation intensities (Fig. 5B) also confirms that the ESM is only really accurate for precipitation events up to 40 mm/d, after which the respective frequencies diverge. The latitudinal and longitudinal mean profiles (Fig. 5C and Fig. 5D) indicate the presence of regional biases.
@@
Our framework demonstrates comparable skill to the QM-based benchmark in correcting the latitude and longitude profiles, for which QM is near optimal by construction (Fig. 5C and Fig. 5D). Comparing the histograms (Fig. 5B and Fig. S3) shows a similar performance of our diffusion model compared to the benchmark while even slightly outperforming it for extreme values.
@@
Our method hence accurately preserves the large-scale precipitation content, while successfully correcting small-scale structure of the precipitation fields, as well as statistical biases in terms of histograms and latitude / longitude profiles (Fig. 5).
@@ 3 Discussion
We introduced a framework based on generative machine learning that enables to do both bias correction and downscaling of Earth system model fields with a single diffusion model. We achieve this by first mapping observational fields and ESM data to a shared embedding space and then applying the learned inverse of the observation embedding transformation to the embedded ESM fields. We learn the inverse transformation with a conditional diffusion model. Although the underlying observational and ESM fields are unpaired, our framework allows for training on paired data, and therefore any supervised machine learning method can be adopted to the task, which allows for more flexibility. The diffusion model is trained on individual samples and has successfully learned to reproduce the statistics of observational data. For the observational ground truth, we chose the ERA5 reanalysis, and for the ESM data to be corrected and downscaled, we chose fields from GFDL-ESM4.
@@
Our diffusion model corrects small-scale biases of the ESM fields, while completely preserving the large-scale structures, which is key for impact assessments, especially with regard to extremes and local impacts in terms of floods or landslides. The diffusion model performs particularly well for extreme events where traditional methods struggle. The method improves the temporal precipitation distribution at a grid cell level and surpasses the state-of-the-art approach (quantile mapping) at correcting spatial patterns. The downscaling performance has also been shown to be excellent. The diffusion model manages to generate small-scale details for the low resolution ESM data, that match those of high resolution observations.
@@
Our method is not specific to ERA5 and GFDL, because the training of the diffusion model does not directly depend on the ESM choice. A specific ESM choice will only modify a hyperparameter in the embedding transformations $f$ and $g$. This, however, requires almost no fine-tuning, as the temporal frequencies can always be matched with quantile mapping. The only parameter that might change for different datasets and use cases is the amount of noise that is added to the observational and ESM datasets. We choose the amount of noise such that the PSDs of the observational ground truth and the ESM fields align beyond a certain scale. This means that we have complete flexibility in deciding which patterns we want to preserve and which we want to correct. This is a major advantage over existing GAN based approaches.
@@
We can decrease the level of detail that is preserved by the diffusion model through increasing the amount of noise added in the transformations $f$ and $g$. This leads to higher uncertainties and therefore increased variance in the outputs of the diffusion model. It is, however, not guaranteed that this increase in variance corresponds to the uncertainties of the ESM. The amount of noise added is directly proportional to the freedom the diffusion model has in generating diverse outputs and inversely proportional to the models’ ability to preserve large-scale patterns.
@@
The downscaled and bias corrected fields will automatically inherit time consistency between different samples up to the noising scale. This means that ESM fields showing two successive days will still look like two successive days after the correction.
@@
We focused on precipitation data over the South American continent, because of its heavily tailed distribution and the pronounced spatial intermittency. Especially at small scales, precipitation data is extremely challenging to model and therefore serves as a reasonable choice to show the frameworks capabilities in a particularly difficult setting. Regional data is chosen due to computational constraints, yet the diverse terrain of our study region, encompassing land, sea, and a wide range of altitudes, enables robust testing of the downscaling and bias correction performance, also given the significant biases of the GFDL model in this region. The extension to global scales is straightforward and requires no changes in the architecture. We intend to include more variables in a consistent manner and on a global scale in future research. Optimizing the inference strategy, with speedup techniques like distillation [17], to decrease the sampling time will prove helpful in this context. Taking the growing number of diffusion model variants into consideration, a comparison between different approaches can help establish the best method suited for the general task of bias correcting and downscaling ESM fields.
@@
It is straightforward to extend our methodology to the downscaling and bias correction of numerical or data driven weather predictions on short to medium-range or even seasonal temporal-scales. This would not require any fundamental changes in the architecture. One would, however, require a target dataset with sufficiently high resolution. The ability of the diffusion model to not disturb temporal consistency between samples can be useful in this scenario.
Declarations
Funding
Funded under the Excellence Strategy of the Federal Government and the Länder through the TUM Innovation Network EarthCare. This is ClimTip contribution #X; the ClimTip project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101137601. PH, SB, and NB acknowledge funding by the Volkswagen Foundation. BP acknowledges funding by the National Key R&D Program of China (2021YFA0718000). YH acknowledges Alexander von Humboldt Foundation for the Humboldt Research Fellowship.
Data Availability
Code Availability
The code will be made available on GitHub.
Competing interests
The authors declare no competing interests.
Ethics approval and consent to participate
Not applicable
Consent for publication
Not applicable
4 Methods

### v2

@@ 2.2 Evaluation of downscaling and bias correction performance
Figure 3: Comparative visualization of individual randomly selected samples. Each row presents three samples of the same dataset. The top row shows GFDL ESM4 data, bilinearly upsampled to 0.25° to match the other fields. The second row shows QM-corrected and the third row diffusion model-corrected GFDL fields. The bottom row shows samples of the original ERA5 data, which are unpaired to the GFDL fields above. Visual inspection shows that the diffusion model correction greatly improves upon the QM correction in terms of producing realistic spatial patterns, since the QM-corrected fields remain way too blurry compared to the HR ERA5 data. The overall large-scale patterns are preserved by the DM. There is no visual difference between the details and sharpness of diffusion model-corrected GFDL fields compared to ERA5.
@@
Figure 4: Comparison of climatologies and model biases. The first row shows the climatology of (A) the diffusion model-corrected GFDL at 0.25°, (B) the GFDL ESM4 model, upsampled to 0.25° and (C) the 0.25° ERA5 data. The second row shows the bias of the GFDL and the QM- and diffusion model-corrections, defined as the difference between long-term temporal averages of all validation samples. Specifically, the temporally averaged bias fields with respect to ERA5 are shown for (D) the diffusion model correction, (E) the uncorrected GFDL and (F) the QM correction. Results indicate a substantial improvement of our diffusion model (A) and the benchmark (C) over just upsampling GFDL to 0.25°. The absolute bias on top of each panel is given by the mean absolute value of the differences over the spatial and temporal dimension with respect to ERA5.
@@
Figure 5: Evaluation of our diffusion model’s performance for downscaling and bias correction. Comparison of GFDL (bilinearly upsampled to 0.25°) (orange) and ERA5 (black) to diffusion model-corrected GFDL (magenta) and QM-corrected GFDL fields (blue) as our benchmark. The Power spectral density (PSD) plot (A) shows that the diffusion model corrects the small-scale spatial details far better than our benchmark. The spectrum aligns very well with the high resolution ERA5 target data. The histograms (B) as well as the latitude (C) and longitude (D) profiles show substantial improvements compared to the uncorrected GFDL data.
We investigate the inference performance of our diffusion model on embedded GFDL data $g(GFDL)$. We compare the downscaling and bias correction performance of our diffusion model to a benchmark consisting of first applying bilinear upsampling followed by QM for bias correction.
Figure 3 presents a qualitative comparison between the different individual precipitation fields. The upsampled $GFDL$ fields, as well as our benchmark are visually too smooth. They therefore appear blurry compared to the ERA5 precipitation fields despite having the same spatial resolution of 0.25°. Our diffusion model produces high-resolution detailed outputs that are visually indistinguishable from the $ERA5$ reanalysis that we treat as the ground truth. We also compared our diffusion model to a different state-of-the-art diffusion model implementation, EDM [24]. The EDM model was trained for the same number of epochs, while taking twice as long for one. The EDM almost perfectly corrects the spectrum (fig. S5A). However in both the histogram (fig. S5B) as well as in latitudinal and longitudinal profiles (fig. S5C and fig. S5D) the EDM model is inferior to our proposed diffusion model. We also compared our method against a VQ-VAE-based generative model, finding that our model outperforms it across these metrics (for details, see SI Sec. SC and fig. S6).
To further validate our choice of architecture, we also compare the diffusion model’s performance against two other state-of-the-art deep learning models, a UNet and a Transformer, using the same experimental setup. The results (fig. S7) show a significant advantage for the DM in reproducing small-scale spatial patterns, by aligning better with the ERA5 reference spectrum (fig. S7A). In contrast, all three models perform comparably well in correcting the overall precipitation distribution and the latitudinal/longitudinal mean profiles (fig. S7B-D). The generative process of the diffusion model is particularly well-suited for correcting the high-frequency spatial details. Another advantage over both deterministic models is the DM’s stochasticity, which allows for the generation of ensembles to quantify uncertainty.
The analysis of temporally averaged precipitation fields shows that the climatology of the diffusion model-corrected GFDL data (Fig. 4A) and the high-resolution ERA5 data (Fig. 4C) is more accurate and less smooth than the climatology of the GFDL data (Fig. 4B). A comparison between the absolute temporally and absolute spatial-temporally averaged diffusion model corrected GFDL and ERA5 fields (Fig. 4D) yields a bias of 0.32 mm d-1. This is a substantial improvement over the original GFDL dataset, which yields a bias of 0.69 mm d-1 (Fig. 4E). Our diffusion model performs comparably with the state-of-the-art bias correction performance of our benchmark, which is by design optimal for this task, at 0.26 mm d-1 (Fig. 4F). For a quantitative comparison including Root Mean Square Error (RMSE) and Pearson correlation for these climatologies, see Table  S1.
@@
There are large differences between the GFDL and ERA5 data in small-scale patterns (Fig. 5A). The histogram of precipitation intensities (Fig. 5B) also confirms that the ESM is only really accurate for precipitation events up to 40 mm d-1, after which the respective frequencies diverge. The latitudinal and longitudinal mean profiles (Fig. 5C and Fig. 5D) indicate the presence of regional biases.
@@
Our framework demonstrates comparable skill to the QM-based benchmark in correcting the latitude and longitude profiles, for which QM is near optimal by construction (Fig. 5C and Fig. 5D). Comparing the histograms (Fig. 5B and fig. S8) shows that our diffusion model is superior compared to the benchmark, strongly outperforming it for extreme values, in particular.
@@
To verify that large-scale patterns are preserved by the diffusion model, we compute image similarity metrics between the low-pass-filtered embedded GFDL data and the low-pass filtered output of the diffusion model. The comparison yields an average structural similarity index value (SSIM [43]) of 0.77 and a Pearson correlation coefficient of 0.90, verifying that large-scale patterns are well preserved by the diffusion model.
We also assess our model’s performance on extreme precipitation events. For this, we use the R95p metric, which is defined as the total annual precipitation from wet days (PR >> 1 mm d-1) that exceed the 95th percentile of our reference period. The difference between the R95p values for the ERA5 and DM corrected GFDL (fig. S9A), the ERA5 and QM corrected GFDL (fig. S9B) and ERA5 and GFDL (fig. S9C), demonstrate that the diffusion model effectively corrects the bias in extreme precipitation events, performing at least as well as the quantile mapping correction. To further test the model’s performance on correcting characteristics of rainfall events in the tail of the distribution, we conduct a return-level analysis for extreme rainfall events (fig. S10). We calculated the average return periods for both moderately extreme (>>50 mm d-1) and very extreme (>>80 mm d-1) events. The raw GFDL model has a significant wet bias, substantially underestimating the return periods (3.33 years and 4.60 years) compared to the ERA5 reference (4.11 years and 7.38 years). Our DM successfully mitigates this bias, yielding more realistic return periods of 4.18 and 7.98 years.
We show that the spatial correlation between the climatologies is improved through our method by computing the Pearson correlation between the temporally averaged fields. The Pearson correlation between ERA5 and GFDL climatology is 0.83, while the correlation between ERA5 and DM-corrected GFDL is 0.98, which is the same as that for the QM-corrected GFDL data. We also investigate how our DM captures the statistics of consecutive dry days (CDD) and consecutive wet days (CWD) compared to the QM benchmark and the raw GFDL (fig. S11). Our diffusion model produces superior CDD (fig. S12A and fig. S12B) and CWD (fig. S12D and fig. S12E) statistics compared to our QM benchmark and GFDL, as shown in the difference plots of CDD / CWD.
@@
Our method therefore accurately preserves the large-scale precipitation content, while successfully correcting small-scale structure of the precipitation fields, as well as statistical biases in histograms and latitude / longitude profiles (Fig. 5). Finally, we confirmed the temporal consistency of our model by analyzing autocorrelation (fig. S13) and seasonal spell duration (fig. S14). We further validated the robustness of our metrics over an extended validation period (1995-2014) (fig. S15).
We also test our framework on a different region of similar size over South Asia. We choose the same GFDL dataset and keep the experimental setup and evaluation identical to the South American region. The setup for quantile mapping the South Asia GFDL data and creating the benchmark data is also the same. We retrained our DM on mapping embedded ERA5 data (over South Asia) to the original ERA5 data. The noising scale in this experiment is the same as for South America, as the PSDs for both regions diverge around the same spatial scale. The evaluation (fig. S16) confirms that our DM successfully corrects precipitation biases in this new region and most notably outperforms the QM baseline in representing small-scale spatial features.
To further assess our framework’s robustness, we conducted an additional experiment using a different ESM. We replaced the GFDL dataset with the MPI-ESM-HR model while keeping the experimental setup and evaluation protocol identical. The MPI and GFDL data diverge at a similar spatial scale in the PSD over the South American domain, allowing us to use the same noising scale hyperparameter s. Quantile delta mapping was applied in the same way as for the GFDL data. Consequently, our diffusion model did not require retraining and could be applied directly to the embedded MPI data at inference. Evaluation on our main metrics (fig. S17) demonstrates our framework’s ability to generalize to different ESMs. Our DM not only restores spatial variability across all scales significantly better than the QM benchmark (fig. S17A), but also shows superior ability to reproduce the frequency of extreme precipitation events (fig. S17B).
2.3 Evaluation of ensemble spread
One of the key strengths of our method lies in its capability to generate a diverse ensemble of downscaled and bias-corrected fields from a single condition. We therefore evaluate the ability of our diffusion model to represent and produce accurate estimates of uncertainty, a critical aspect for robust climate modeling and decision-making. We generate a 50-member DM ensemble by running the model 50 times, each conditioned on the same low-resolution ERA5 year, producing one-year trajectories. The corresponding high-resolution year serves as the ground truth. Our results demonstrate that the DM ensemble effectively reproduces the correct precipitation patterns, as shown by the close alignment between the ensemble mean and the high resolution ground truth of ERA5 over the annual cycle (fig. S18). Probabilistic performance, evaluated using CRPS, highlights that the DM significantly outperforms a bilinear baseline, with lower mean CRPS values (0.76 mm d-1 vs 0.90 mm d-1), as well as better temporally and spatially averaged CRPS (fig. S19). Furthermore, we confirm that the DM ensemble produces well-calibrated uncertainty estimates with a spread-skill plot. Our model achieves near-perfect alignment with the 1:1 line, indicating an accurate representation of uncertainty (fig. S20). For more details see SI Sec. SD.1.
2.4 Evaluation on future climate scenarios
Evaluating the performance of downscaling models is crucial for their application in climate impact studies under future climate scenarios. We assess our diffusion model’s ability to preserve climate change signals in the underlying ESM simulations by applying it to a high-emission future scenario (SSP5-8.5). Figure 6 compares the relative climate change signal between the late 21st century (2081-2100) and the historical period (1995–2014) for annual mean and annual extreme precipitation. We find that our downscaled 0.25° fields successfully capture the mean precipitation change, closely matching the pattern and magnitude shown in the original 1° GFDL data (Fig. 6A and Fig. 6B). The diffusion model also robustly preserves the climate change signal for extreme precipitation indices, including Rx1Day (wettest day for each year) and R95p (Fig. 6C - Fig. 6F). The spatial patterns of change for the extremes are well-reproduced in the DM-corrected output compared to the original model data. Notably, slight differences are observed in the northwestern domain (Fig. 6C and Fig. 6E), where the DM-correction projects a slightly stronger increase in extreme events under SSP5-8.5. A slight increase in extremes aligns with the diffusion model’s bias correction capabilities, reflecting its role in addressing the known under-representation of extreme precipitation in the original GFDL simulations.
Furthermore, we demonstrate that our conditionally trained diffusion model generalizes robustly to unseen future emission scenarios by accurately preserving regional precipitation trends without requiring retraining. We analyze the full annual mean precipitation timeseries from 2015 to 2100 over two representative regions, one exhibiting a strong negative trend and one with a pronounced positive trend (fig. S21). For each region, we compare the annual mean precipitation from the original GFDL SSP5-8.5 data at 1° with the DM-corrected output at 0.25° resolution. The diffusion model consistently preserves the direction and magnitude of the trends found in the original GFDL data across the entire timeseries, for both the negative (fig. S21 blue) and positive trend (fig. S21 red) regions. This demonstrates the model’s ability to maintain physically meaningful long-term changes in precipitation, further supporting its generalization capability to future scenarios. Note that the absolute values do not have to coincide, as our model corrects the bias and hence the numerical values. Our model can generalize to unseen climates, preserving the trends, since there is no decrease in performance during inference on GFDL SSP5-8.5 data. Note that our set-up generalized to unseen climate scenarios without any external constraints. The reason why our model preserves trends well is likely given by the fact that the trend is dominated by the large-scale patterns and our model learned to rely on the large-scale patterns of the condition and only generates small-scale patterns.
Figure 6: Comparison of relative climate change signals. We compute the relative climate change signal between the late 21st century (2081-2100) under the GFDL SSP5-8.5 scenario and the historical GFDL period (1995–2014). In (A) and (B), we show that our diffusion model successfully preserves the mean precipitation climate change signal in the downscaled 0.25° GFDL fields, matching the change of the original 1° GFDL data. Positive values (red) indicate an intensification of precipitation, while negative values (blue) indicate reductions. In (C-D) and (E-F) we evaluate how well the DM-correction preserves the climate change signal for extreme events in historical and future scenarios. Both the Rx1Day (C-D) as well as R95p in (E-F) show that the DM-downscaling does preserve the climate change signal for extreme events. There are only slight differences over the north western part of our fields, where the DM-correction predicts more slightly more extremes for the SSP5-8.5 scenario. This is in line with the bias correction capabilities of the DM, correcting the under-representation of extreme precipitation in the original GFDL data.
@@ 3 Discussion
We introduced a framework based on generative machine learning that allows both bias correction and downscaling of Earth system model fields with a single diffusion model. We achieve this by first mapping observational fields and ESM data to a shared embedding space and then applying the learned inverse of the observation embedding transformation to the embedded ESM fields. We learn the inverse transformation with a conditional diffusion model. Although the underlying observational and ESM fields are unpaired, our framework allows for training on paired data (between observations and embedded observations, see above) and therefore any supervised machine learning method can be adopted to the task, which allows for more flexibility. Supervised methods are often superior in performance and more natural for the downscaling application. The diffusion model is trained on individual samples and has successfully learned to reproduce the statistics of observational data. For the observational ground truth, we chose the ERA5 reanalysis, and for the ESM data to be corrected and downscaled, we chose fields from GFDL-ESM4.
We demonstrated our framework’s robustness and generalizability in two additional experiments (Sec. 2.2). When applying the model to a new geographical region in South Asia with the same ESM, the DM requires retraining to adapt to the new regional characteristics. In contrast, when applying the framework to an entirely different ESM (MPI-ESM) over the South American region, the core DM did not need to be retrained since the same noising scale hyperparameter could be used. For different ESMs a new noising scale hyperparameter could be necessary, requiring retraining of our DM with a different noising scale; however this depends on the choice of the spatial scale below which bias correction is desired, and for comparable outputs, we recommend to keep the noising scale $s$ fixed for different ESMs. For example, to correct multiple ESMs at once, one can use the most heavily biased model to select the noising scale. A single diffusion model can then be trained to correct all ESMs at once, saving significant computational resources during inference. In general, we expect that many ESMs (like the MPI and GFDL model we use) will have similar spatial scales up to which they can capture realistic spatial precipitation features, because they have a similar resolution and have similar limitations from parameterization schemes. In all cases, readjusting the computationally inexpensive Quantile Delta Mapping (QDM) is a required step in the embedding process. The results will also depend on the specific quantile mapping scheme, QDM is chosen to preserve trends.
@@
Our diffusion model corrects small-scale biases of the ESM fields, while completely preserving the large-scale structures, which is key for impact assessments, especially with regard to extremes and local impacts in terms of floods or landslides. The diffusion model performs particularly well for extreme events where traditional methods struggle. The method improves the temporal precipitation distribution at the grid cell level and surpasses the state-of-the-art approach (quantile mapping) in correcting spatial patterns. The downscaling performance has also been shown to be excellent. The diffusion model manages to generate small-scale details for the low resolution ESM data, that match those of high resolution observations. Our model preserves relevant information from the large scales, such as trends and extremes, and generates bias corrected and downscaled precipitation fields with adequate uncertainties.
We show that our method is robust in the out-of-distribution setting of downscaling and bias-correcting the SSP5-8.5 future emission scenario. It is critical for impact assessments that our model is able to accurately preserve the climate change signal of the original SSP5-8.5 data.
A key innovation of our approach is the embedding strategy, which makes the training process independent of the source ESM (apart from a single data-dependent hyperparameter setting the spatial scale below which the fields are corrected), which not only allows the framework to be flexibly applied to downscale and bias-correct a wide range of ESMs but also allows it to be used with different state-of-the-art machine learning backbone models. Another key advantage of our framework is its data efficiency. In our conditional approach the model only needs to learn how to generate small-scale features given the large-scale ones. The task is considerably less demanding than that of unconditional models (e.g., Hess et al. [19]), which must learn the entire data distribution from scratch during training. This data efficiency makes our method applicable to datasets with shorter record lengths than ERA5, such as newer observational products.
Indeed, comparing results for generated climatologies between our conditional DM and the unconditional consistency model (CM) by Hess et al. [19], it becomes apparent that the CM struggles to learn the target distribution accurately, leading to blurring (fig. S2) that would hinder applications for impact assessments.
@@
Our method is not specific to ERA5 and GFDL because the training of the diffusion model does not directly depend on the ESM choice. A specific ESM choice will only modify a hyperparameter in the embedding transformations $f$ and $g$. This, however, requires almost no fine-tuning, as the temporal frequencies can always be matched with quantile mapping. The only parameter that might change for different datasets and use cases is the amount of noise that is added to the observational and ESM datasets. We choose the amount of noise such that the PSDs of the observational ground truth and the ESM fields align beyond a certain scale. This means that we have complete flexibility in deciding which patterns we want to preserve and which we want to correct. This is a major advantage over existing GAN based approaches.
@@
We can decrease the level of detail that is preserved by the diffusion model through increasing the amount of noise added in the transformations $f$ and $g$. The amount of noise added is directly proportional to the freedom the diffusion model has in generating diverse outputs and inversely proportional to the model’s ability to preserve large-scale patterns.
@@
The downscaled and bias corrected fields will automatically inherit time consistency between different samples up to the noising scale. This means that ESM fields showing two successive days will still look like two successive days after the correction. Future work could build a video diffusion model that inputs and outputs full time series instead of single frames, in order to guarantee time consistency across all scales.
@@
We focused on precipitation data over the South American continent, because of its heavily tailed distribution and the pronounced spatial intermittency. Especially at small scales, precipitation data is extremely challenging to model and therefore serves as a reasonable choice to show the framework’s capabilities in a particularly difficult setting. Regional data is chosen due to computational constraints, yet the diverse terrain of our study region, encompassing land, sea, and a wide range of altitudes, enables robust testing of the downscaling and bias correction performance, also given the substantial biases of the GFDL model in this region. We also conducted additional experiments for another region over South Asia, and using another ESM, namely the MPI-ESM-HR, in order to confirm the generality of our approach. The extension to global scales is straightforward and requires no major changes in the architecture. We intend to include more variables in a consistent manner on a global scale in future research. Optimizing the inference strategy, with speedup techniques such as distillation [27], to decrease the sampling time will prove helpful in this context.
As for any ML model, the ability to generate the rarest extremes is limited by their frequency in the training data. Our conditional approach helps mitigate this to some extent by inheriting the large-scale patterns for these events directly from the ESM.
@@
It is straightforward to extend our methodology to downscaling and bias correction of numerical or data-driven weather predictions on short- to medium-range or even seasonal temporal scales. This would not require any fundamental changes to the architecture. This would, however, require a target dataset with sufficiently high resolution. The ability of the diffusion model to not disturb the temporal consistency between samples can be useful in this scenario. Future work could then focus on extending this model to a multivariate setting, which would be essential for weather prediction and for assessing physical consistency between variables.
4 Materials and Methods